TAYLOR IS 7 YEARS OLDER THAN ANTHONY.
A)WRITE A EQUATION FOR THIS SITUATION
B) SOLVE THE PROBLEM BY WORKING BACKWARDS IF TAYLOR WAS 13 YEARS OLD

Answer by TimothyLamb(4379)

x = y + 7
13 = y + 7
y = 13 - 7
y = 6
---
anthony is 6
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php